OpenStudy (anonymous):

Use the Binomial Theorem to expand the binomial. (d – 4b)3

OpenStudy (anonymous):

Can you solve this one? \[\sum_{k=0}^{3} \left(\begin{matrix}3 \\ k\end{matrix}\right) d ^{3-k}(-4b) ^{k}\]

OpenStudy (anonymous):

no im not good at math

OpenStudy (anonymous):

Do you know what mean by \[\sum_{k=0}^{3}" number"\]

OpenStudy (anonymous):

The expantion of the first thing i wrote is \[=\left(\begin{matrix}3 \\ 0\end{matrix}\right) d ^{3}+\left(\begin{matrix}3 \\ 1\end{matrix}\right)d ^{2}(-4b)^{1}+\left(\begin{matrix}3 \\ 2\end{matrix}\right)d ^{1}(-4b)^{2}+\left(\begin{matrix}3 \\ 3\end{matrix}\right)(-4b)^{3}\] If you don't get the idea of that you should probably go back a few pages in your book and reveiw it